OpenStudy (anonymous):

Calculus This is the last of the questions I didn't understand on my worksheet, please please help? the function s(t) describes the motion of a particle moving along a line. For each function, (a) find the velocity function of the particle at any time t >_ 0, (b) identify the time interval(s) when the particle is moving in a positive direction, (c) identify the time interval(s) when the particle is moving in a negative direction, and (d) identify the time(s) when the particle changes its direction s(t)=t^2-7t+10

OpenStudy (mertsj):

The velocity function is the first derivative of the distance.

OpenStudy (loser66):

why are you stuck? and where?

OpenStudy (anonymous):

I don't understand

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

The position function is \(s(t)\). The rate of change of position is the derivative of \(s(t)\), or \(s'(t)\). The "rate of change of position" is the definition of "velocity," so we call \(s'(t)=v(t)\). To find the velocity function, it's only a matter of differentiating \(s(t)\) with respect to \(t\). \[s(t)=t^2-7t+10~~\Rightarrow~~s'(t)=v(t)=2t-7\]

OpenStudy (anonymous):

thank you:)

OpenStudy (anonymous):

and to find b do I have to graph it?

OpenStudy (kainui):

Nope, but if you don't immediately know it would be wise to graph it so you start to develop an intuition of what a graph might look like and where it is moving in a positive direction so that later you can do it without graphing it and just sort of vaguely visualize it in your head.

OpenStudy (anonymous):

How can I solve it without graphing?

OpenStudy (kainui):

Think about it. This is something you need to figure out on your own because it's a simple concept that you are 100% capable of seeing. Ask yourself: when is the particle moving in the positive direction?

OpenStudy (kainui):

After you graph it you will be able to figure out on your own clearly how to solve it without graphing it. Think about the slope. Derive your own understanding.

OpenStudy (anonymous):

@Loser66 please help me with this last question, please

OpenStudy (loser66):

what I can help? for a) velocity is derivative of displacement, so , just take s'(t)

OpenStudy (anonymous):

"Moving in a positive direction" seems to refer to "when \(s(t)\) is increasing." Likewise, "moving in a negative direction" sounds like "when \(s(t)\) is decreasing." Apply the first derivative test.

OpenStudy (anonymous):

Also, the "times when the particle changes direction," assuming what I said earlier is right, means "find the local extrema," i.e. when \(s(t)\) stops increasing/decreasing and beings to decrease/increase, respectively.

OpenStudy (loser66):

I think so, too. That's why I ask him "why are you stuck" . He worked on it many times yesterday.

OpenStudy (anonymous):

aaah ok.

OpenStudy (anonymous):

I wasn't sure

OpenStudy (anonymous):

thank you so much